OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer? 114 116 112 118

OpenStudy (anonymous):

2x+(2x+2)=226 x = 56 2x + 2=114

OpenStudy (anonymous):

Even integers are the integers which are divisible by 2 and has a gap or difference of 2 like 2, 4, 6 and so on.. So, if one even integer we let x, then the other even integer will be: (x + 2). x + x + 2 = 226 Or, 2x = 224 Now can you solve it for x???

OpenStudy (jiteshmeghwal9):

Let the two consecutive integers be (x)&(x+2).Then, (x)+(x+2)=226 =>2x+2=226 =>2x=226-2 =>2x=224 =>x=224/2=112. So, larger integer=(x+2)=112+2=114.

OpenStudy (anonymous):

Just check the solution of Jiteshmeghwal9 and no need to solve on your own...

OpenStudy (anonymous):

thanks